# Generalized electrical Lie algebras

### v3

@@ Abstract.
We generalize the electrical Lie algebras originally introduced by Lam and Pylyavskyy in several ways. To each Kac-Moody Lie algebra $\mathfrak{g}$ we associate two types (vertex type and edge type) of the generalized electrical algebras. The electrical Lie algebras of vertex type are always subalgebras of $\mathfrak{g}$ and are flat deformations of the nilpotent Lie subalgebra of $\mathfrak{g}$. In many cases including $sl_{n}$, $so_{n}$, and $sp_{2n}$ we find new (edge) models for our generalized electrical Lie algebras of vertex type. Finding an edge model in general is an interesting an open problem.
@@ 1. Introduction and main results
It turns out that these relations are defining. This was conjectured for finite types in [8, Conjecture 5.2], which which was partially confirmed in [10] and [3, 4].
@@
In the notation of Theorem 1.2 the relations (1.2) provide a presentation of $\mathfrak{g}^{({\bf a})}$. Moreover, $\mathfrak{g}^{({\bf a})}$ is a flat deformation of the nilpotent part ${\mathfrak{n}}=<e_{i},i\in I>$ of $\mathfrak{g}$ in the sense that $\mathfrak{g}^{({\bf a})}\cong\mathfrak{n}$ for any ${\bf a}\in\mathbb{C}^{I}$ as a naturally filtered Lie algebra.
We prove Theorems 1.3 in Section 3.2 (along with their generalization to any coefficient algebra $R$ containing all $a_{i}$) by utilizing a mini-theory of flat deformations of associative and Lie algebras which we present for the reader’s convenience in Appendix.
Note that if we view all $a_{i}$ as formal parameters (after replacing $\mathbb{C}$ with $\mathbb{C}[{\bf a}]=\mathbb{C}[a_{i},i\in I]$), then $\mathbb{C}[{\bf a}]\otimes\mathfrak{g}$ is graded by $\mathbb{Z}^{I}_{\geq 0}$ via $\deg a_{i}=\deg e_{i}=-\deg f_{i}$, $i\in I$. In particular, $u_{i}$ is homogeneous with $\deg u_{i}=\deg a_{i}$ and both $\mathfrak{g}^{({\bf a})}$ and $U(\mathfrak{g}^{({\bf a})})$ are $\mathbb{Z}^{I}_{\geq 0}$-graded.
Now we construct a real form $\mathfrak{g}^{({\bf b})}$ for the electrical Lie algebra $sl_{n}^{({\bf b})}$.
@@
Remark 1.5.
This realization with all $a_{i}=\sqrt{-1}$ (i.e., all $b_{j}=1$) was first discovered by T. Lam and A. Postnikov as it is mentioned in [7].
The realization with all $a_{i}$ imaginary is a “real form” of the electric Lie algebra of type $sl_{n}$. We produce more real forms for $\mathfrak{g}^{(\bf a)}$ below and in Section 2.
@@
for all distinct $i,j\in I$ where $b_{ij}=0$ unless $a_{ij}=-1$. We refer to $\mathfrak{g}^{({\bf b})}$ as the generalized electrical Lie algebra of the edge type $\mathfrak{g}$ associated with the Lie algebra $\mathfrak{g}$ (so it is logical to think of $\mathfrak{g}^{({\bf a})}$ as the vertex type). Ultimately, on Theorems 1.4 and 1.6, we say that an edge model of an electric Lie algebra of type $\mathfrak{g}$ is any subalgebra of $\mathfrak{g}$ isomorphic to $\mathfrak{g}^{({\bf b})}$.
@@
$\Omega_{6}^{(\bf b)}=\begin{pmatrix}0&b_{1}b_{2}b_{3}b_{4}&0&0&0&0\\ -b_{1}b_{2}b_{3}b_{4}&0&-b_{2}b_{3}b_{4}&0&0&0\\ 0&b_{2}b_{3}b_{4}&0&b_{3}b_{4}&0&0\\ 0&0&-b_{3}b_{4}&0&-b_{4}&0\\ 0&0&0&b_{4}&0&1\\ 0&0&0&0&-1&0\end{pmatrix}.$
@@
Our next result also gives an edge model for dihedral, type $D$, and affine $A$ types, however, unlike Theorem 1.10, we do not expect any intertwiner from the vertex model.
@@
(a) Let $\mathfrak{g}=\mathfrak{g}_{A}$ be a Kac-Moody algebra of rank $2$ with $I=\{1,2\}$ such that $a_{21}\leq-2$. Then the assignments $u_{1}\mapsto e_{1}$, $u_{2}\mapsto e_{2}+b_{12}f_{1}$, define an injective homomorphism of Lie algebras $\mathfrak{g}^{(\bf b)}\hookrightarrow\mathfrak{g}$.
$b:=a_{21}a_{1}a_{2}$
@@
(b) If all $b_{i}\neq 0$ then, under then there is an isomorphism of Lie algebras $J\cong sl_{n+1}^{({\bf b^{\prime}})}$ for some choice of the parameters ${\bf b^{\prime}}$ (which we specify in the proof) and the Lie algebra $sp^{({\bf b})}_{2n}$ is naturally isomorphic to $sl_{n}^{(b_{1},\ldots,b_{n-2})}\ltimes sl_{n+1}^{({\bf b^{\prime}})}$.
@@
Remark 1.14.
It is well-known that $\mathfrak{h}\ltimes\mathfrak{g}\cong\mathfrak{h}\oplus\mathfrak{g}$ for any Lie algebra $\mathfrak{g}$ and it subalgebra $\mathfrak{h}$ under the right adjoint action of $\mathfrak{h}$ on $\mathfrak{g}$ (namely, the diagonal copy of $\mathfrak{h}$ in $\mathfrak{h}\ltimes\mathfrak{g}\cong$ commutes with $(0,\mathfrak{g})$).
Moreover, applying this to $\mathfrak{g}=sp_{n}$, $\mathfrak{h}=sp_{n-1}$, Theorems 1.8 and 1.13(b) produce, after an appropriate localization, an isomorphism
recovering [10, Thm. 2.3.1] [^1] The electric algebras of $B_{n}$ type in [10] correspond to $C_{n}$ type in our approach..
We conclude with an illustration of Theorem 1.13 for $n=3$.
@@
The Chevalley generators of $sp_{6}$ are given in terms of the folding of those for $sl_{6}$ under the flip $i\mapsto 6-i$:
@@
$\tilde{u}_{4}=e_{4}+b_{1}f_{5},\,\tilde{u}_{5}=e_{5}$
the generators of ${sp}^{(\bf b)}_{6}$ are
$u_{1}=\tilde{u}_{1}+\tilde{u}_{5},\,u_{2}=\tilde{u}_{2}+\tilde{u}_{4},\,u_{3}=\tilde{u}_{3}.$
Introduce the elements of ${sp}^{(\bf b)}_{6}$
$w_{3}=u_{3},\,w_{2}=[u_{2},[u_{2},u_{3}]],\,w_{1}=[u_{1},[u_{1},w_{2}]].$
These elements on the one hand generate the ideal $J$ and on the other hand a copy of $sl_{4}^{({\bf b^{\prime}})}$ where ${\bf b^{\prime}}=\{-32b^{2}_{1}b^{2}_{2}\,,-8b^{2}_{2}\}$. This shows a splitting of $sp_{6}^{({\bf b})}$ into a semidirect product as claimed in the statement of the theorem.
Moreover, as Remark 1.14 claims there is copy of $sl_{3}^{({\bf b^{\prime}})\backslash-8b^{2}_{2}}$ that commutes with the above copy of $sl_{4}^{({\bf b^{\prime}})}$. It is generated by the elements
@@ 2. Other vertex and edge models of electrical Lie algebras
In this section, we generalize Theorem 1.4, Example 1.9 and Theorems 1.10, 1.11 to other semisimple and Kac-Moody algebras.
@@
$u_{i}\mapsto e_{0}-a_{0}\delta_{i,0}({\bf f}-\frac{a_{0}}{2}[{\bf f},[{\bf f},f_{0}]]-a_{0}\sum\limits_{k\geq 3}\frac{1}{k!}(ad~{}{\bf f}+a_{0}[f_{0},{\bf f}])^{k-1}([{\bf f},[{\bf f},f_{0}]]))\ ,$
@@ 3. Proof of main results
Thus, all conditions of Proposition 4.1 are met. Thus, the proposition guarantees that the relations (1.2) give a presentation of $B=U(\mathfrak{g}^{(\bf a)})$ as an associative algebra over $R$ hence of the Lie algebra $\mathfrak{g}^{(\bf a)}$ over $R$.
@@
Since $sl_{n}^{(b)}$ also annihilates $v^{1}\in V$, $sl_{n}^{(b)}$ identifies with a Lie subalgebra of $sp_{n-1}$. Then the dimension argument once again implies that $sl_{n}^{({\bf b})}\cong sp_{n-1}$ as well for even $n$.
@@
Let $\hat{\hat{R}}=\mathbb{C}[a_{i},i\in I]$ an let $\hat{R}$ be its subalgebra generated by all $b_{ij}:=a_{ji}a_{i}a_{j}$ whenever $a_{ij}=-1$ in the assumptions of Theorems 1.10 and 2.3. Clearly, $\hat{R}=\mathbb{C}[b_{ij}]$.
Indeed, let $u_{i}=u_{i}^{\bf b}\in\hat{R}\otimes\mathfrak{g}$, $i\in I$ be the elements as in (the right hand sides of) Theorems 2.3 and 1.10. Let $v_{ij}^{\bf b}\in\hat{R}\otimes\mathfrak{g}$ be the left hand side of (1.3) in each case.
@@
Thus, all conditions of Proposition 4.1 are met. Thus, the proposition guarantees that the relations of Theorems 1.10 and 2.3 give a presentation of $B=U(\mathfrak{g}^{(\bf b)})$ as an associative algebra over $R$ hence, of the Lie algebra $\mathfrak{g}^{(\bf b)}$ over $R$.
Theorems 1.6, 2.3, and 1.10 are proved. $\square$
@@
Recall that ${sp}_{2n}$ contains $sl_{n}$ and splits as a $sl_{n}$ module
@@
and
$u^{\prime}_{n-k}=Ad^{2}_{u_{n-k}}\circ Ad^{2}_{u_{n-k+1}}\circ\dots\circ Ad^{2}_{u_{n-1}}(u_{n})$
There is an isomorphism of Lie algebras $\phi:sl_{n+1}^{(\bf b^{\prime})}\rightarrow J$ such that $\phi(v_{i})=u^{\prime}_{i}$, where $v_{i}$ are the generators of $sl_{n+1}^{({\bf b^{\prime}})}$ and ${\bf b^{\prime}}=(b^{\prime}_{1},\ldots,b^{\prime}_{n-1})$
$b^{\prime}_{k}=-2^{2(n-k)+1}\prod_{i=k}^{n-1}b^{2}_{i}$
$k=1,\ldots n-1$.
Moreover $[u^{\prime}_{n-k},u^{\prime}_{n-k\pm 1}]=(-2)^{k}\left(\prod_{i=1}^{k-1}b_{n-i}\right)[u_{n-k},u^{\prime}_{n-k\pm 1}]$ for $11\leq i\leq n-$ 1 and hence
$[u^{\prime}_{n-k}+(-2)^{k+1}\left(\prod_{i=1}^{k}b_{n-i}\right)u_{n-k},u^{\prime}_{n-k\pm 1}]=0$
Observe finally, that the elements
@@
form a copy of $sl_{n}^{(b_{1},\ldots,b_{n-2})}$ after the appropriate localization.
@@
Therefore, we obtain
$(ad~{}u_{i})^{3}(u_{j})=(ad~{}e_{i})^{3}(e_{j})+t_{k}^{3}t_{\ell}(ad~{}f_{k})^{3}(f_{\ell})$
because $[e_{i},f_{kk\ell}]=0$, $[f_{k},e_{iij}]=0$.

### v5

@@ Abstract.
We generalize the electrical Lie algebras originally introduced by Lam and Pylyavskyy in several ways. To each Kac-Moody Lie algebra $\mathfrak{g}$ we associate two types (vertex type and edge type) of the generalized electrical algebras. The electrical Lie algebras of vertex type are always subalgebras of $\mathfrak{g}$ and are flat deformations of the nilpotent Lie subalgebra of $\mathfrak{g}$. In many cases including $sl_{n}$, $so_{n}$, and $sp_{2n}$ we find new (edge) models for our generalized electrical Lie algebras of vertex type. Finding an edge model in general is an interesting open problem.
@@ 1. Introduction and main results
It turns out that these relations are defining. This was conjectured for finite types in [8, Conjecture 5.2], which was partially confirmed in [10] and [3, 4].
@@
In the notation of Theorem 1.2 the relations (1.2) provide a presentation of $\mathfrak{g}^{({\bf a})}$. Moreover, $\mathfrak{g}^{({\bf a})}$ is a flat deformation of the nilpotent part ${\mathfrak{n}}=<e_{i},i\in I>$ of $\mathfrak{g}$ in the sense that $\mathfrak{g}^{({\bf a})}\cong\mathfrak{n}$ for any ${\bf a}\in\mathbb{C}^{I}$ as naturally filtered vector spaces.
We prove Theorem 1.3 in Section 3.2 (along with its generalization to any coefficient algebra $R$ containing all $a_{i}$) by utilizing a mini-theory of flat deformations of associative and Lie algebras which we present for the reader’s convenience in Appendix.
Note that if we view all $a_{i}$ as formal parameters (after replacing $\mathbb{C}$ with $\mathbb{C}[{\bf a}]=\mathbb{C}[a_{i},i\in I]$), then $\mathbb{C}[{\bf a}]\otimes\mathfrak{g}$ is graded by the root lattice $Q:=\bigoplus\limits_{i\in I}\mathbb{Z}\alpha_{i}$ via $\deg a_{i}=\deg e_{i}=-\deg f_{i}=\alpha_{i}$, $i\in I$. In particular, $u_{i}$ is homogeneous with $\deg u_{i}=\alpha_{i}$ and both $\mathfrak{g}^{({\bf a})}$ and $U(\mathfrak{g}^{({\bf a})})$ are $Q$-graded.
Now we construct a “real form” $\mathfrak{g}^{({\bf b})}$ of the electrical Lie algebra $sl_{n}^{({\bf b})}$.
@@
A bit paradoxically, if all $a_{i}$ are imaginary hence $b_{j}$ in Theorem 1.4. We produce more real forms of $\mathfrak{g}^{(\bf a)}$ below and in Section 2.
@@
for all distinct $i,j\in I$ where $b_{ij}=0$ unless $a_{ij}=-1$. We refer to $\mathfrak{g}^{({\bf b})}$ as the generalized electrical Lie algebra of the edge type $\mathfrak{g}$ associated with the Lie algebra $\mathfrak{g}$ (so it is logical to think of $\mathfrak{g}^{({\bf a})}$ as the vertex type). Ultimately, on Theorems 1.4 and 1.5, we say that an edge model of an electric Lie algebra of type $\mathfrak{g}$ is any subalgebra of $\mathfrak{g}$ isomorphic to $\mathfrak{g}^{({\bf b})}$.
@@
$\Omega_{6}^{(\bf b)}=\tiny{\begin{pmatrix}0&b_{1}b_{2}b_{3}b_{4}&0&0&0&0\\ -b_{1}b_{2}b_{3}b_{4}&0&-b_{2}b_{3}b_{4}&0&0&0\\ 0&b_{2}b_{3}b_{4}&0&b_{3}b_{4}&0&0\\ 0&0&-b_{3}b_{4}&0&-b_{4}&0\\ 0&0&0&b_{4}&0&1\\ 0&0&0&0&-1&0\end{pmatrix}.}$
@@
Our next result also gives an edge model for dihedral, type $D$, and affine $A$ type, however, unlike Theorem 1.9, we do not expect any intertwiner from the vertex model.
@@
(a) Let $\mathfrak{g}=\mathfrak{g}_{A}$ be a Kac-Moody algebra of rank $2$ with $I=\{1,2\}$ such that $a_{21}\leq-2$. Let $b_{12}:=a_{21}a_{1}a_{2}$. Then the assignments $u_{1}\mapsto e_{1}$, $u_{2}\mapsto e_{2}+b_{12}f_{1}$, define an injective homomorphism of Lie algebras $\mathfrak{g}^{(\bf b)}\hookrightarrow\mathfrak{g}$.
@@
(b) If all $b_{i}\neq 0$, then there is an isomorphism of Lie algebras $J\cong sl_{n+1}^{({\bf b^{\prime}})}$ for some choice of the parameters ${\bf b^{\prime}}$ (which we specify in the proof) and the Lie algebra $sp^{({\bf b})}_{2n}$ is naturally isomorphic to $sl_{n}^{(b_{1},\ldots,b_{n-2})}\ltimes sl_{n+1}^{({\bf b^{\prime}})}$.
Remark 1.13.
It is well-known that $\mathfrak{h}\ltimes\mathfrak{g}\cong\mathfrak{h}\oplus\mathfrak{g}$ for any Lie algebra $\mathfrak{g}$ and its subalgebra $\mathfrak{h}$ under the right adjoint action of $\mathfrak{h}$ on $\mathfrak{g}$ (namely, the diagonal copy of $\mathfrak{h}$ in $\mathfrak{h}\ltimes\mathfrak{g}$ commutes with $(0,\mathfrak{g})$).
Moreover, applying this to $\mathfrak{g}=sp_{n}$, $\mathfrak{h}=sp_{n-1}$, Theorems 1.7 and 1.12(b) produce, after an appropriate localization, an isomorphism
recovering [10, Theorem 2.3.1] [^1] The electric algebras of type $B_{n}$ in [10] correspond to the type $C_{n}$ in our approach..
@@
The Chevalley generators of $sp_{6}$ are given in terms of the folding of the standard generators for $sl_{6}$ under the flip $i\mapsto 6-i$:
@@
$\tilde{u}_{4}=e_{4}+b_{1}f_{5},\,\tilde{u}_{5}=e_{5}\ .$
The generators of ${sp}^{(\bf b)}_{6}$ are $u_{1}=\tilde{u}_{1}+\tilde{u}_{5},\,u_{2}=\tilde{u}_{2}+\tilde{u}_{4},\,u_{3}=\tilde{u}_{3}$. Introduce the elements of ${sp}^{(\bf b)}_{6}$: $w_{3}=u_{3},\,w_{2}=[u_{2},[u_{2},u_{3}]],\,w_{1}=[u_{1},[u_{1},w_{2}]]$. These elements, on the one hand, generate the ideal $J$ and, on the other hand, a copy of $sl_{4}^{({\bf b^{\prime}})}$ where ${\bf b^{\prime}}=\{-32b^{2}_{1}b^{2}_{2}\,,-8b^{2}_{2}\}$. This shows a splitting of $sp_{6}^{({\bf b})}$ into a semidirect product as in Theorem 1.12(c).
Moreover, as Remark 1.13 claims there is a copy of $sl_{3}^{({\bf b^{\prime}})\backslash-8b^{2}_{2}}$ that commutes with the above copy of $sl_{4}^{({\bf b^{\prime}})}$. It is generated by the elements
@@ 2. Other vertex and edge models of electrical Lie algebras
In this section, we generalize Theorem 1.4, Example 1.8 and Theorems 1.9, 1.10 to other semisimple and Kac-Moody algebras.
@@
$u_{i}\mapsto e_{i}-a_{0}\delta_{i,0}({\bf f}-\frac{a_{0}}{2}[{\bf f},[{\bf f},f_{0}]]-a_{0}\sum\limits_{k\geq 3}\frac{1}{k!}(ad~{}{\bf f}+a_{0}[f_{0},{\bf f}])^{k-1}([{\bf f},[{\bf f},f_{0}]]))\ ,$
@@ 3. Proof of main results
Thus, all conditions of Proposition 4.1 are met. The proposition guarantees that the relations (1.2) give a presentation of $B=U(\mathfrak{g}^{(\bf a)})$ as an associative algebra over $R$ hence of the Lie algebra $\mathfrak{g}^{(\bf a)}$ over $R$.
@@
Since $sl_{n}^{(b)}$ also annihilates $v^{1}\in V$, $sl_{n}^{(b)}$ is a Lie subalgebra of $sp_{n-1}$. Then the dimension argument once again implies that $sl_{n}^{({\bf b})}\cong sp_{n-1}$ as well for even $n$.
@@
Let $\hat{\hat{R}}=\mathbb{C}[a_{i},i\in I]$ and let $\hat{R}$ be its subalgebra generated by all $b_{ij}:=a_{ji}a_{i}a_{j}$ whenever $a_{ij}=-1$ in the assumptions of Theorems 1.9 and 2.3. Clearly, $\hat{R}=\mathbb{C}[b_{ij}]$.
Indeed, let $u_{i}=u_{i}^{\bf b}\in\hat{R}\otimes\mathfrak{g}$, $i\in I$ be the elements as in (the right hand sides of) Theorems 2.3 and 1.9. Let $v_{ij}^{\bf b}\in\hat{R}\otimes\mathfrak{g}$ be the left hand side of (1.3) in each case.
@@
Thus, all conditions of Proposition 4.1 are met. The proposition guarantees that the relations of Theorems 1.9 and 2.3 give a presentation of $B=U(\mathfrak{g}^{(\bf b)})$ as an associative algebra over $R$ hence, of the Lie algebra $\mathfrak{g}^{(\bf b)}$ over $R$.
Theorems 1.5, 2.3, and 1.9 are proved. $\square$
@@
Recall that ${sp}_{2n}$ contains $sl_{n}$ and splits as a $sl_{n}$-module
@@
and $u^{\prime}_{n-k}=Ad^{2}_{u_{n-k}}\circ Ad^{2}_{u_{n-k+1}}\circ\dots\circ Ad^{2}_{u_{n-1}}(u_{n})$. There is an isomorphism of Lie algebras $\phi:sl_{n+1}^{(\bf b^{\prime})}\rightarrow J$ such that $\phi(v_{i})=u^{\prime}_{i}$, where $v_{i}$ are the generators of $sl_{n+1}^{({\bf b^{\prime}})}$, ${\bf b^{\prime}}=(b^{\prime}_{1},\ldots,b^{\prime}_{n-1})$ where $b^{\prime}_{k}=-2^{2(n-k)+1}\prod\limits_{i=k}^{n-1}b^{2}_{i}$ $k=1,\ldots n-1$.
Moreover $[u^{\prime}_{n-k},u^{\prime}_{n-k\pm 1}]=(-2)^{k}\left(\prod\limits_{i=1}^{k-1}b_{n-i}\right)[u_{n-k},u^{\prime}_{n-k\pm 1}]$ for $1\leq i<n$ hence $[u^{\prime}_{n-k}+(-2)^{k+1}\left(\prod\limits_{i=1}^{k}b_{n-i}\right)u_{n-k},u^{\prime}_{n-k\pm 1}]=0$ Observe finally, that the elements
@@
generate a copy of $sl_{n}^{(b_{1},\ldots,b_{n-2})}$ after the appropriate localization.
@@
Therefore, we obtain $(ad~{}u_{i})^{3}(u_{j})=(ad~{}e_{i})^{3}(e_{j})+t_{k}^{3}t_{\ell}(ad~{}f_{k})^{3}(f_{\ell})$ because $[e_{i},f_{kk\ell}]=0$, $[f_{k},e_{iij}]=0$.